# Video Is Worth a Thousand Images: Exploring the Latest Trends in Long Video Generation

### v1

@@ Abstract.
An image may convey a thousand words, but a video—composed of hundreds or thousands of image frames—tells a more intricate story. Despite significant progress in multimodal large language models (MLLMs), generating extended videos remains a formidable challenge. As of this writing, OpenAI’s Sora [110], the current state-of-the-art system, is still limited to producing videos of up to one minute in length. This limitation stems from the complexity of long video generation, which requires more than generative AI techniques for approximating density functions—essential aspects such as planning, story development, and maintaining spatial and temporal consistency present additional hurdles. Integrating generative AI with a divide-and-conquer approach could improve scalability for longer videos while offering greater control. In this survey, we examine the current landscape of long video generation, covering foundational techniques like GANs and diffusion models, video generation strategies, large-scale training datasets, quality metrics for evaluating long videos, and future research areas to address the limitations of the existing video generation capabilities. We believe it would serve as a comprehensive foundation, offering extensive information to guide future advancements and research in the field of long video generation.
@@
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
A transformative breakthrough in image generation accompanied the success of LLMs. DALL-E 2 [79] represented a significant leap beyond traditional Generative Adversarial Network (GAN) and Variational Autoencoder (VAE) models, thanks to its ability to interpret natural language and render a wide range of concepts and styles. It excels at creating photo-realistic images from both realistic and imaginative prompts. Other leading generative AI systems, such as Stable Diffusion 3 [19] and MidJourney [106], also showcase remarkable capabilities in producing photo-realistic visuals based on real and imagined ideas.
Generating videos is significantly more challenging than creating text or images. Unlike static images or text, videos are dynamic, featuring evolving camera angles, motion, deformations, and occlusion patterns. Producing realistic, long-duration videos requires continuously generating new content while preserving consistency. Videos are composed of multiple frames, with their semantic content evolving. Here, a frame’s ”semantic content” refers to a conceptual map that represents how various elements within the frame—such as objects, actions, and interactions—contribute to the overall understanding of the video. Fig. 1 illustrates an example of a single-scene video featuring a girl dancing with repetitive motions against a static background. This example highlights minimal changes in semantic content across frames, with no introduction of new characters or actions. Such a video clip may be classified as a single scene whose semantic content remains unchanged or broadly consistent throughout. In contrast, Fig. 2 shows an example of a multi-scene video where the number of objects changes and a new actor appears at a later frame. In this case, the semantic content of the frames changes as the video progresses.
@@
Due to the complexities of dynamic scenes, early video generation models were limited to producing short clips lasting only a few seconds, often animating a single static frame without incorporating varying backgrounds or objects. For example, Make-A-Video  [88] and RunwayML Gen-2  [104] generate 4-5 second videos using a single animated frame with little change in semantic content. CogVideo  [35] is among the first long video generation models to create extended videos using autoregressive transformers. However, it operated based on a single prompt and also exhibited minimal changes in the semantic content of the video. Phenaki  [120], employing autoregressive video transformers, is one of the first models to generate long videos with dynamic semantic content based on multiple prompts. Similarly, Gen-L-Video  [124] video uses a diffusion model to combine short video clips into a single, continuous video. Recently, Sora  [110] has established a new state-of-the-art in video generation. It utilizes diffusion transformers and sampling from a compressed spatiotemporal representation and produces videos that are visually impressive and rich in semantic content. Another state of art model in long video generation is Gen-3 Alpha by RunwayML  [105]. It is a commercial model that can generate photo-realistic videos up to 10 seconds long and is also based on diffusion transformers. Fig. 3 shows the evolution of these long video generation techniques, which are still in their early stages compared to human-made commercial videos, especially regarding video length and narrative organization.
Figure 3. Evolution of long video generation models. Later models like SORA [110] and Gen-3 Alpha [105] focus more on the quality of videos than the length of the video.
Figure 4. Most papers focusing on long video generation were published in 2023 and 2024.
@@ 1.1. Survey Contributions — Need for Summarizing Long Video Generation Methods
Long video generation presents numerous challenges beyond crafting and maintaining consistent storylines across scenes. These include the lack of large-scale video datasets with detailed captions and the requirement of significant computational resources. Despite these hurdles, long video generation has emerged as a new frontier in generative AI, offering potential applications in entertainment, education, medicine, marketing, and gaming. This has attracted significant research interest, leading to a rapid increase in studies. As shown in Fig. 4, most papers on long video generation have been published within the last two years.
@@
Owing to these interests and opportunities, the time is right to summarize the state-of-the-art in long video generation and discuss the associated challenges, progress, and future directions to support the advancement of the field. To our knowledge, there are only two related long video-generation surveys  [50] and  [154]. The former [50] delves into the latest trends in long video generation, emphasizing the divide-and-conquer and autoregressive approaches as two primary themes. It also examines photo-realism trends and generative paradigms, such as VAE, GAN, and diffusion-based models. However, while it highlights the divide-and-conquer approach, which simplifies the complexity of long videos by breaking them into smaller, manageable chunks, a detailed exploration of this methodology—such as how short videos can be seamlessly integrated into longer narratives—is lacking. Our work aims to bridge this gap by thoroughly analyzing the various dimensions of the divide-and-conquer strategy and its role in addressing the challenges of long video generation. In contrast, the latter [154] offers a broader overview of the video generation field, covering topics like long videos, video editing, super-resolution, datasets, and metrics. However, long video generation is only one of many topics discussed. It highlights the need for a focused, in-depth study on long video generation. Our contribution addresses this need by comprehensively analyzing this emerging field and highlighting key approaches, challenges, and future prospects.
@@
The proposed work addresses gaps in existing literature by incorporating recent studies and conducting a comprehensive analysis of the divide-and-conquer approach. Specifically, we focus on underexplored aspects, such as agent-based networks and methods for transitioning from short to long videos, which are missing from current reviews. Furthermore, we extend our examination to methodologies and papers that fall outside the divide-and-conquer and autoregressive categories and the latest research not covered by previous surveys. In addition, we dive deeply into the literature, emphasizing the algorithms, models, and input control techniques employed in generating long videos, offering a more holistic and detailed perspective on the field.
@@ 1.2. Survey Focus — Techniques, Challenges & Key Questions in Long Video Generation
Video generation utilizes various techniques, such as sampling from latent space [57], creating small video segments or images, generating intermediate frames (”divide and conquer”)[3.2], employing autoregressive methods[3.1] to predict future frames based on initial ones, and enhancing latent state representations for longer videos. Training video generation models presents challenges due to the higher computational requirements and more extensive memory needs for video datasets. Many video generation models build on pre-trained image models[ [37][62],  [5]], enhancing attention mechanisms to ensure consistency between adjacent frames, as video is essentially a sequence of frames. Some long video models are developed by extending short video generation models[ [147][11][124]] and improving control mechanisms for longer content. Another significant aspect of video generation is input guidance[[4.1],[4.3],[4.2] ]. Unlike image or short video generation, guidance is crucial for long video generation and is often based on text embedding like CLIP [76]. LLMs [3.2.1] are crucial in guiding the video generation process by understanding world dynamics, object trajectories, and action groupings, using their extensive knowledge to inform the creation of videos. When evaluating the quality of the generated videos[[6.1],[6.2],[ 6.3],[ 6.4]], it is important to assess the quality of individual frames, the fluidity of motion, and the overall aesthetic appeal. It is essential to ensure that the generated videos align with the input text and maintain the consistency of entities throughout the video, such as cars and actors. This has sparked interest in investigating innovative research directions in the field and addressing the following relevant research questions:
@@
    How can we manage resource requirements for training and inference with long videos?
  4. (4)
    How can we handle control signals (conditioning) for extended video sequences?
@@ 1.3. Survey Methodology
For this survey, we conducted searches across several conferences, including but not limited to ICCV, ECCV, CVPR, ICLR, ICML, IEEE, TPAMI, IEEE Neural Networks, AAAI, WACV, ICIP, NeurIPS, KDD, ACCV, IJCV, AI Open, and CVIU. We used keywords such as ”video generation,” ”long video generation,” and ”LLM-guided video generation.” Additionally, we searched academic databases, including Google Scholar, IEEE Xplore, ACM Transactions, and Scopus, focusing on the term ”long video generation” for our survey.
Our survey covers papers published between 2021 and 2024 (till August end), with a focus on ”video generation” and ”generative AI”. We gathered 100+ articles through snowball sampling, using keywords like text-to-video, generative AI, visual interpretation, and extended video generation.
@@ 1.4. Survey Organization
We will begin by discussing the foundational framework for video generation, including embedding and LLMs, to set the stage for more advanced topics. The goal is to familiarize readers with these fundamental components, enabling them to explore these building blocks according to their level of expertise. Next, we will explore backbone mechanisms for video generation, like divide and conquer autoregressive and global enhancements. We will explore input guidance mechanisms, including strategies like LLM guidance, and categorize them into different levels based on the depth of control the LLM exerts. We will also address the necessary image and video diffusion model modifications to facilitate such control. We will also discuss the post-processing pipelines required to achieve high temporal and spatial quality in videos generated by diffusion models. We will then discuss datasets used to train video generation models. We will then talk about metrics used to measure generated video quality. Finally, we will talk about future trends and open challenges.
@@ 2. Long Video Generation: Backbone Architectures and Methods
Progress in long video generation builds on advancements in many foundational building blocks. These include GANs-based architecture 2.1, autoencoders  2.2, Transformers-based models 2.3, LLMs and language understanding 2.4 and image and video diffusion models 2.5.
@@ 2.1. GAN Based Video Generation
Generative Adversarial Networks (GANs) were considered the state-of-the-art (SOTA) for generative tasks from around 2014 to the early 2020s, but recent advancements such as diffusion models and transformer-based methods have surpassed them in both performance and versatility, particularly in applications like image synthesis and video generation. This section explores the formulation of GANs and highlights the pivotal papers that have shaped advancements in the field. GANs [22] are inspired by game theory and consist of two core components: the Generator and the Discriminator. The Generator’s purpose is to convert random noise, typically sampled from a simple, uniform distribution, into data samples, which can range from images to videos. Simultaneously, the Discriminator assesses these samples, classifying them as real (originating from the training dataset) or fake (created by the Generator). This GAN formulation is the basis of image and video generation, which further sections explain.
@@ 2.1.1. GAN Based Image Generation
GANs initially revolutionized image generation, dominating the field. Here, we examine key ideas and milestones in GAN literature, organized by timeline.
Early GANS: GANs  [22] was the first to generate images using adversarial networks but employed simple feedforward neural networks for both the Generator and the Discriminator. DCGANs  [77] extends the GAN architecture by incorporating convolutional layers, making them more suitable for image data. They generated images with a resolution of 64×64. LAPGAN  [12] increased the resolution of images by generating them multi-scale. It consists of multiple GANs, each generating images at different resolutions. GANs, DCGANs, and LAPGAN are primarily designed to generate images based on random noise vectors. These models don’t take text guidance, and text-based GANs add text-based control, which will be discussed in the next section.
GAN Text Input:  [149] is a multistage text-to-image GAN that can create high-quality images. It has two GAN stacks on each other. The first one takes the text and generates a low-resolution image. The second one takes both text and input images and creates high-quality images. AttnGAN  [138] uses attention mechanisms to create images from text, allowing it to focus on specific words or phrases in the input description.
Style/Image Transfer: GAN was the first generative model to generate high-quality artistic images, and one of the key innovations was transferring an artistic style like Vincent Van Gogh’s to real-world pictures and image translation. CycleGAN  [157] does image-to-image translation and consists of two generators and two discriminators. StyleGAN  [43] primarily focuses on generating high-quality, diverse images, particularly faces. It introduced disentangled latent space. Latent space is where a vector of N dimensions represents each image. Projecting different high-level attributes like skin color, hairstyle, etc., on distinct dimensions in latent space provides excellent editing capabilities for realistic image generation, semantic manipulation, and local editing. StyleGAN opened the doors for high-level image manipulation. StyleGAN2  [119] is an improvement over StyleGAN for better quality images. pix2pix  [40] specifically designed for image-to-image translation tasks. It learns a conditional generative model and generates an output image conditioned on the input image. GAN also revolutionized video generation, which will be explored in  2.1.2.
@@ 2.1.2. Video/Multi Frame Generation
Early Attempts: Early research in video synthesis primarily concentrated on video prediction like in [65], which involves generating future frames based on a sequence of previously observed ones.  [122] extended the convolutional model to include two distinct convolution networks. It decomposes video into a static background and a moving foreground, generating them with 2D and 3D convolutional networks. VGAN can generate 32 frames of realistic videos of golf courses, beaches, train stations, etc. VGAN is an unconditional video generator, i.e., without incorporating any auxiliary input such as text. ’Multi-stage Dynamic Generative Adversarial Networks’ [136] predicts future video frames based on the first frame. It improved existing models by generating 32 frames with 128 x 128 resolution in a time-lapse pattern. The first stage generates a time-lapse video with realistic content in its two-stage model. In contrast, the second stage optimizes the results from the first stage, primarily by incorporating dynamic motion information to enhance realism. The models mentioned above are not prompt-based, and prompt-based models will be explored in the following section.
@@
Long Video Generation Using GAN: DIGAN  [144] can create 128 frame video. It introduced an INR (Implicit Neural Representations) based video generator that improves motion dynamics by manipulating space and time coordinates differently and a motion discriminator that efficiently identifies unnatural motions without observing long frame sequences. StyleGan-V  [90] improved the state of the art and was built on StyleGAN2  [44]. It can generate high-resolution 1024-long videos by designing a holistic discriminator that aggregates temporal information by simply concatenating frames’ features to decrease the training cost.

### v2

@@ Abstract.
An image may convey a thousand words, but a video, composed of hundreds or thousands of image frames, tells a more intricate story. Despite significant progress in multimodal large language models (MLLMs), generating extended videos remains a formidable challenge. As of this writing, OpenAI’s Sora [1], the current state-of-the-art system, is still limited to producing videos of up to one minute in length. This limitation stems from the complexity of long video generation, which requires more than generative AI techniques for approximating density functions. Critical elements, such as planning, narrative construction, and spatiotemporal continuity, pose significant challenges. Integrating generative AI with a divide-and-conquer approach could improve scalability for longer videos while offering greater control. In this survey, we examine the current landscape of long video generation, covering foundational techniques such as GANs and diffusion models, video generation strategies, large-scale training datasets, quality metrics for evaluating long videos, and future research areas to address the limitations of existing video generation capabilities. We believe it would serve as a comprehensive foundation, offering extensive information to guide future advancements and research in the field of long video generation.
@@
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The success of LLMs brought a transformative breakthrough in image generation. DALL-E 2 [7] surpassed traditional GANs and VAEs by interpreting natural language and generating diverse concepts and styles, excelling in photorealistic outputs. Other systems, such as Stable Diffusion 3 [8] and MidJourney [9], also demonstrate strong capabilities in creating realistic visuals.
Video generation is far more complex than text or images due to dynamic elements like motion, occlusion, and evolving semantic content—the conceptual mapping of objects, actions, and interactions. Single-scene videos (e.g., a girl dancing against a static background, Fig. 1) maintain consistent semantics, while multi-scene videos (Fig. 2) introduce new objects or actors, altering semantic content over time.
@@
Due to the complexities of dynamic scenes, early video generation models were limited to producing short clips lasting only a few seconds, often animating a single static frame without incorporating varying backgrounds or objects. For example, Make-A-Video  [12] and RunwayML Gen-2  [13] generate 4-5 second videos using a single animated frame with little change in semantic content. CogVideo  [14] is among the first long video generation models to create extended videos using autoregressive transformers. However, it operated based on a single prompt and also exhibited minimal changes in the semantic content of the video. Phenaki [15], which employs autoregressive video transformers, is one of the first models to generate long videos with dynamic semantic content based on multiple prompts. Similarly, Gen-L-Video  [16] employs a diffusion model to merge short video clips into a seamless, continuous video. Sora  [1] has established a new state-of-the-art in video generation. Sora is also known as the ChatGPT of the video generation world. It utilizes diffusion transformers and sampling from a compressed spatiotemporal representation, producing videos that are visually impressive and rich in semantic content. Another state-of-the-art model in long video generation is Gen-4 Alpha by RunwayML  [17]. It is a commercial model that can generate photo-realistic videos up to 10 seconds long and is also based on diffusion transformers. Fig. 4 illustrates the evolution of these long video generation techniques, which are still in their early stages compared to human-made, commercial videos, particularly in terms of video length and narrative organization.
Figure 3. Most papers focusing on long video generation were published in 2023–2025.
Figure 4. Evolution of long video generation models. Later models, such as SORA [1] and Gen-4 Alpha [17], focus more on video quality than duration.
@@ 1.1. Survey Contributions — Need for Summarizing Long Video Generation Methods
Long video generation presents numerous challenges beyond crafting and maintaining consistent storylines across scenes. These include the lack of large-scale video datasets with detailed captions and the requirement of significant computational resources. Despite these challenges, long video generation has emerged as a transformative frontier in generative AI, unlocking novel possibilities in various fields, including entertainment, education, healthcare, marketing, and gaming. This potential has sparked substantial research attention, and publication rates have accelerated dramatically in recent years. As illustrated in Figure 3, the field has experienced explosive growth, with more than two-thirds of all academic work produced within the past 24 months—a testament to its dynamic transformation.
@@
Given these interests and opportunities, it is time to summarize the state-of-the-art in long video generation and discuss the associated challenges, progress, and future directions that will support the advancement of the field. To our knowledge, there are only two related long video-generation surveys  [18] and  [19]. The former [18] explores the latest trends in long video generation, highlighting the divide-and-conquer and autoregressive approaches as two primary themes. It also examines photo-realism trends and generative paradigms, such as VAE, GAN, and diffusion-based models. However, while it highlights the divide-and-conquer approach, which simplifies the complexity of long videos by breaking them into smaller, manageable chunks, a detailed exploration of this methodology, such as how short videos can be seamlessly integrated into longer narratives, is lacking. Our work aims to bridge this gap by thoroughly analyzing the various dimensions of the divide-and-conquer strategy and its role in addressing the challenges of long video generation. In contrast, the latter [19] provides a broader overview of the video generation field, encompassing topics such as long videos, video editing, super-resolution, datasets, and metrics. However, long-video generation is only one of many topics discussed. It highlights the need for a focused and in-depth study on the generation of long videos. Our contribution addresses this need by providing a comprehensive analysis of this emerging field and highlighting key approaches, challenges, and prospects.
@@
The proposed work addresses the gaps in the existing literature by incorporating recent studies and conducting a comprehensive analysis of the divide-and-conquer approach. Specifically, we focus on underexplored aspects, such as agent-based networks and methods for transitioning from short to long videos, which are missing from current reviews. Furthermore, we extend our examination to methodologies and papers that fall outside the divide-and-conquer and autoregressive categories, as well as the latest research not covered by previous surveys. In addition, we dive deeply into the literature, emphasizing the algorithms, models, and input control techniques employed in generating long videos, offering a more holistic and detailed perspective on the field.
@@ 1.2. Survey Focus — Techniques, Challenges & Key Questions in Long Video Generation
Video generation utilizes various techniques, such as sampling from latent space [20], creating small video segments or images, generating intermediate frames (”divide and conquer”)[3.2], employing autoregressive methods [3.1] to predict future frames based on initial ones, and improving latent state representations for longer videos. Training video generation models presents challenges due to the higher computational requirements and more extensive memory needs for video datasets. Many video generation models are based on pretrained image models [ [21],  [22],  [23]], which enhance attention mechanisms to ensure consistency between adjacent frames, as video is essentially a sequence of frames. Some long video models are developed by extending short video generation models[ [24][25][16]] and improving control mechanisms for longer content. Another significant aspect of video generation is input guidance[[5.1],[5.3],[5.2] ]. Long video generation requires stronger guidance than images or short clips, typically anchored in text embeddings, such as CLIP [26]. Here, large language models (Section 3.2.1) take center stage: they decode physical world dynamics, forecast object interactions, and choreograph multistep actions, leveraging their pre-trained knowledge to steer generation toward coherent, long-form outputs. When evaluating the quality of the generated videos[[7.1],[7.2],[ 7.3],[ 7.4]], it is important to assess the quality of the individual frames, the fluidity of motion and the overall aesthetic appeal. Ensuring that generated videos remain faithful to the input text while preserving entity consistency (e.g., cars, actors) across frames is a critical challenge. This has motivated researchers to explore novel directions in the field, raising key questions that warrant further investigation.
@@
    Discussion of Long-video generation strategies covering segmented stitching, auto-regressive frames, and full latent-space synthesis.
@@ 1.3. Survey Methodology
For this survey, we conducted searches across several conferences, including but not limited to CVPR, ICLR, NeurIPS, IJCV, IEEE-TVCG, TPAMI, AAAI, CVIU, IEEE Neural Networks, ICML, ACM Multimedia, IEEE-WACV. We used keywords such as ”video generation,” ”long video generation,” and ”LLM-guided video generation.” Additionally, we searched academic databases, including arXiv, Google Scholar, IEEE Xplore, ACM Transactions, and Scopus, focusing on the term ”long video generation” for our survey. Our survey covers papers published between 2021 and 2025 (as of May 2025), with a focus on ”video generation” and ”generative AI”. We gathered over 190+ articles through snowball sampling, using keywords such as text-to-video, generative AI, visual interpretation, and extended video generation.
@@ 1.4. Survey Organization
We will begin by discussing the foundational frameworks for video generation, including embedding and LLMs, to set the stage for more advanced topics. The goal is to familiarize readers with these fundamental components, enabling them to explore these building blocks according to their level of expertise and interest. Next, we will explore backbone mechanisms for video generation, including divide-and-conquer autoregressive models and the use of implicit latent spaces. We will then explore Tokenization Strategies. We will explore input guidance mechanisms, including strategies such as LLM guidance, and categorize them into different levels based on the depth of control the LLM exerts over them. We will also address the necessary modifications to the image and video diffusion models to facilitate such control. We will also discuss the post-processing pipelines required to achieve high temporal and spatial quality in videos generated by diffusion models. We will then discuss the datasets used to train video generation models, as outlined in Section datasets . We will then discuss the metrics used to measure generated video quality, as outlined in Section metrics. Finally, we will talk about future trends and open challenges.
@@ 2. Long Video Generation: Backbone Architectures and Methods
Progress in long-video generation builds on advancements in many foundational building blocks. These include GANs-based architecture 2.1, Autoencoders  2.2, Transformers-based models 2.3, LLMs and language understanding 2.4, and Image and Video Diffusion models 2.5.
@@ 2.1. GAN Based Video Generation
GANs [27] dominated generative tasks from 2014 until the early 2020s, although diffusion models and transformer-based approaches in terms of performance and versatility have since surpassed them. The fundamental GAN framework [27] consists of two competing components: a Generator that creates samples from random noise and a Discriminator that evaluates their authenticity. While this adversarial architecture established the foundation for image and video generation, newer methods have advanced beyond its capabilities, as discussed in subsequent sections.
@@ 2.1.1. GAN Based Image Generation
GANs initially revolutionized image generation, dominating the field of image generation. Here, we examine key ideas and milestones in GAN literature, organized by timeline.
Early GANS: GANs  [27] was the first to generate images using adversarial networks but employed simple feedforward neural networks for both the Generator and the Discriminator. DCGANs  [28] extend the GAN architecture by incorporating convolutional layers, making them more suitable for image data. They generated images with a resolution of 64×64. LAPGAN  [29] increased the resolution of images by generating them multi-scale. It consists of multiple GANs, each generating images at different resolutions. GANs, DCGANs, and LAPGAN are primarily designed to create images based on random noise vectors. These models lack text guidance, but text-based GANs incorporate text-based control, which we will discuss in the next section.
GAN Text Input: StackGAN  [30] is a multistage text-to-image GAN that generates high-quality images. It has two GAN stacks stacked on top of each other. The first one takes the text and generates a low-resolution image. The second one takes both text and input images and creates high-quality images. AttnGAN  [31] uses attention mechanisms to create images from text, allowing it to focus on specific words or phrases in the input description.
Style/Image Transfer: StyleGAN [32] was the first generative model to generate high-quality artistic images, and one of the key innovations was transferring an artistic style like Vincent Van Gogh’s to real-world pictures and image translation. CycleGAN  [33] does image-to-image translation and consists of two generators and two discriminators. StyleGAN2  [34] primarily focuses on generating high-quality, diverse images, particularly faces. It introduced disentangled latent space. Latent space is where a vector of N dimensions represents each image. Projecting different high-level attributes, such as skin color and hairstyle, onto distinct dimensions in a latent space provides excellent editing capabilities for realistic image generation, semantic manipulation, and local editing. StyleGAN2[34] opened the doors for high-level image manipulation. StyleGAN2 [34] is an improvement over StyleGAN, producing higher-quality images. pix2pix  [35] specifically designed for image-to-image translation tasks. It learns a conditional generative model and generates an output image conditioned on the input image. GAN also revolutionized video generation, which we will explore in  2.1.2.
@@ 2.1.2. Video/Multi Frame Generation
Early Attempts: [36], which generated future frames from observed sequences.  [37] advanced this by using separate 2D and 3D convolutional networks for static backgrounds and moving foregrounds, producing 32-frame unconditional videos of various scenes. Further development came with  [38]’s two-stage model, which first generated 128×128 resolution time-lapse videos from a single frame and then enhanced them with dynamic motion information. These early works laid the important foundations for unconditional video generation before the advent of prompt-based approaches.
@@
Long Video Generation Using GAN: DIGAN  [40] can create 128 frame video. It introduced an INR (Implicit Neural Representations) based video generator that improves motion dynamics by manipulating space and time coordinates differently and a motion discriminator that efficiently identifies unnatural motions without requiring long frame sequences. StyleGan-V  [41] improved the state of the art and was built on StyleGAN2  [34]. It can generate high-resolution 1024-long videos by designing a holistic discriminator that aggregates temporal information by simply concatenating frame features, thereby decreasing the training cost.